### v2

@@ Abstract.
The increasing affordability of robot hardware is accelerating the integration of robots into everyday activities. However, training robots to automate tasks typically requires physical robots and expensive demonstration data from trained human annotators. Consequently, only those with access to physical robots produce demonstrations to train robots. To mitigate this issue, we introduce EVE, an iOS app that enables everyday users to train robots using intuitive augmented reality visualizations without needing a physical robot. With EVE, users can collect demonstrations by specifying waypoints with their hands, visually inspecting the environment for obstacles, modifying existing waypoints, and verifying collected trajectories. In a user study ($N=14$, $D=30$) consisting of three common tabletop tasks, EVE outperformed three state-of-the-art interfaces in success rate and was comparable to kinesthetic teaching—physically moving a real robot—in completion time, usability, motion intent communication, enjoyment, and preference ($mean_{p}=0.30$). We conclude by enumerating limitations and design considerations for future AR-based demonstration collection systems for robotics.
augmented reality, demonstration collection, imitation learning
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; , 2024; ccs: Human-centered computing Mixed / augmented realityccs: Computer systems organization Robotics
@@ 1. Introduction
Figure 1. EVE enables everyday users to collect data to train a real robot using intuitive augmented reality (AR) visualizations without needing a physical robot. Using our interface, the user can move the AR robot by setting waypoints with the hand, visually inspect the real world to avoid obstacles, modify the collected trajectory by reverting to a previous waypoint, and verify the collected trajectory by replaying the task with the AR robot. See supplementary materials for the videos of real-world deployment.
@@
The decreasing costs of robot components and systems are accelerating the integration of robots into everyday life. With companies offering robotic arms at a few hundred dollars, portable, lightweight, and mobile robots are being trained to assist individuals with a wide range of everyday tasks (Elephant Robotics, 2023; Robot, 2023; Shaw et al., 2023a). However, training robots to automate tasks often requires expensive manually curated human demonstrations. Annotators are hired and trained to kinesthetically move a robot’s joints and gripper to a sequence of configurations to accomplish a task. For example, to train a robot to clean a table, annotators move the gripper towards a cloth, close the gripper to grasp the cloth, and then begin moving it around to wipe the table.
@@
As such, there are no mechanisms for everyday users, who are considering purchasing a robot, to train robots without first purchasing an expensive robot and learning to move bulky machinery to produce effective demonstrations. Consequently, only a limited number of roboticists and developers can determine the range of robotic applications.
@@
Recently, a prototype, called AR2-D2, demonstrated the possibility of training robots without the need for a physical robot (Duan et al., 2023). AR2-D2 enables users to capture videos of themselves manipulating objects in various environments using smartphones or tablets. Furthermore, it has proven that the quality of data gathered through AR visualization is comparable with data collected using traditional approaches with real robots. However, AR2-D2 has only been evaluated on simple, short-horizon tasks such as pressing a computer mouse or picking up a plastic bowl. Ecologically valid user tasks, such as cleaning a table, require longer interactions between the robot and its environment. Additionally, the system’s usability has not been analyzed, which is crucial for understanding the usability and utility for every user.
We introduce EVE, an iOS app that allows everyday users to train robots with human-centric AR visualizations and features, even if they do not have access to a robot. To collect demonstrations, users can specify waypoints for the trajectory with their fingers or make the AR robot track the hand in real time. During collection, users can visually inspect the real world to avoid obstacles, modify the collected trajectory by reverting to a previous waypoint, and verify the collected trajectory by replaying the task with the AR robot.
@@
To understand the opportunities and challenges of AR-based demonstration collection, we conducted a formative study with 10 participants with varying experience in robotics, using AR2-D2 as the baseline system and three new AR visualizations. We identified six usability challenges users faced and implemented seven additional features to improve EVE’s accuracy, flexibility, and verifiability.
@@
We conducted a user study with 14 participants to compare EVE with state-of-the-art demonstration collection interfaces on three typical tabletop tasks. EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in task completion time, system usability, motion intent communication, user enjoyment, and overall preference ($mean_{p}=0.30$). Participants customized their workflows with multiple visualizations and features to collect demonstrations for complex, long-horizon tasks.
In summary, our contributions include:
  • EVE, an iOS app that enables everyday users to train robots using intuitive AR visualizations and features
  • Formative study that identifies opportunities and challenges for improving the usability of various demonstration collection interfaces
  • User study that demonstrates how people use diverse collection interfaces
  • A discussion on designing future AR-based demonstration collection interfaces
@@ 2. Related Work
2.1. Human Demonstrations Collection for Robotics
Imitation learning enables robots to effectively automate tasks by learning from human demonstrations (Zare et al., 2023; Wang et al., 2023; Shridhar et al., 2022). A common approach for collecting demonstration data includes kinesthetic teaching, where a human physically guides the robot through a desired trajectory (Argall et al., 2009). Despite the direct haptic feedback, this mechanism can be slow and tedious (Osentoski et al., 2010). Another popular collection interface is teleoperation, where the user controls the robot remotely using specialized controllers, such as 3D space mouse (Zhu et al., 2022), VR controllers(Ebert et al., 2021; Brohan et al., 2022; Seo et al., 2023), smartphones(Mandlekar et al., 2018), and haptic devices(Shaw et al., 2023b; Toedtheide et al., 2023), which can be costly and challenging to use due to high latency and lack of intuitiveness (Akgün and Subramanian, 2011). More recent efforts also involve capturing human expert demonstrations with physical robots using specialized hardware like ALOHA(Zhao et al., 2023), GELLO(Wu et al., 2023), DexCap(Wang et al., 2024), and UMI(Chi et al., 2024). However, these studies often prioritize data collection efficiency and model training with only a few exceptions, including minimum studies (Kapadia et al., 2017; Szafir and Szafir, 2021) on examining user preference and intuitiveness in their systems.
@@ 2.1. Human Demonstrations Collection for Robotics
Prior works compared kinesthetic teaching and teleoperation on simple tasks such as manipulating cups (Fischer et al., 2016; Akgün and Subramanian, 2011). The common findings are that kinesthetic teaching is preferred over teleoperation based on collection time, ease of use, and accuracy of demonstrations. In contrast to this work, our paper compared the performance of traditional collection interfaces as well as state-of-the-art AR-based collection interfaces. Furthermore, we focus on three realistic tabletop tasks with varying levels of difficulty, leading to a more comprehensive comparison between different collection interfaces.
Most closely related to our work is AR2-D2, a system for collecting demonstrations that do not require skilled demonstrators and real robots and resituating demonstration collection outside the lab setting (Duan et al., 2023). It is an iOS app that people can use to project AR robots into the real world and collect waypoint information with their hands. The user study demonstrates that AR2-D2 is intuitive to use and requires no training in contrast to four other common demonstration collection interfaces. However, AR2-D2 did not experiment with different user interface designs and only offered one way to control and visualize the robot. In addition, it only allows the collection of waypoint data, which is not useful for training the robot to do dynamic tasks such as tossing a ball. In contrast, EVE focuses on the human-centric aspect of the demonstration collection process by providing three new AR visualizations and seven new features derived from the formative study. Furthermore, EVE enables the collection of both waypoint data and trajectory data, allowing users to select suitable collection methods based on task types.
@@ 2.2. Augmented Reality and Robotics
Traditional approaches to improving human-robot interaction, such as through robots’ movements (Dragan et al., 2013; Mainprice et al., 2010), gestural motion (Huang and Mutlu, 2013; Chidambaram et al., 2012), and gaze outputs (Mutlu et al., 2006; Kalegina et al., 2018), are limited due to the challenge of presenting expressive physical feedback that goes beyond internal capabilities. In contrast, AR provides visual feedback in one’s line of sight, opening up exciting new opportunities for human-robot interaction research. A recent taxonomy of augmented reality and robotics has identified five purposes and benefits of visual augmentation (Suzuki et al., 2022). We focus on analyzing works related to motion intent communication since our goal is to improve the usability of the system by making it easier for the user to understand the goal of the AR robot.
@@
AR interfaces improve a robot’s motion intent communication through spatial information. Prior works have experimented with different designs based on factors such as path, waypoints, velocity, arrival times, etc. (Walker et al., 2018; Zollmann et al., 2014; Rosen et al., 2017). Most relevant to our work is Rosen et al., who visualized the robotic arm’s motion intent by overlaying the expected robot poses with a mixed reality head-mounted display. They evaluated the performance of the visualization against 2D displays and no visualizations and found that the system allows a human to more quickly (-62% in time) and accurately (+16% in accuracy) tell where the robot is going to move. In addition to visualizing the robot’s motion intent, they iterated on the system to allow the robot to replan an intended trajectory with MoveIt motion planner (mov, 2023) with the same start and end points, with the final trajectory chosen by the human. This has been shown to improve task completion time, precision, accuracy, usability, and task load compared to the baseline systems. Unlike this work, our application is in the form of an iOS app, and the AR visualization is applied with a phone or a tablet. This increases the potential to enable more people to collect demonstrations for training robots compared to wearable displays. Furthermore, instead of projecting the entire pose of the robot in AR, we move the AR robot directly to the goal pose due to the lack of safety concerns. Similar to how they allow the user to replan the path, we offer a redo function that allows users to reselect the trajectory they desire.
3. EVE Prototype 1
@@
Figure 2. Prototype 1 of EVE includes three new AR visualizations. AR kinesthetic teaching enables the robot to track the user’s hand in real-time. Path history displays trajectory along collected waypoints. Invisible robot allows users to toggle the robot body off and only view a cylinder that shows the robot’s end effector position and the gripper state (green = open, red = close).
@@
While prior work has compared common demonstration collection interfaces and various AR visualizations, there is a lack of research investigating user preferences for AR-based demonstration collection systems. Since AR2-D2 only provides one way to control and visualize the robot’s movement, we developed the first EVE prototype with three new AR visualizations for the formative user study: AR kinesthetic teaching, path history, and invisible robot (Figure 2). Below, we describe the design and implementation of the new AR visualizations.
@@ 3.1. AR Kinesthetic Teaching (AR-KT)
Prior work suggests that kinesthetic teaching outperforms teleoperation in system usability, collection time, and task success rate (Duan et al., 2023). This is primarily due to the real-time feedback and intuitiveness when manually guiding the robot with the hand. Inspired by these findings, we implemented AR-based kinesthetic teaching (AR-KT). When AR-KT is activated, the AR robot tracks the movement of the user’s hand and adjusts the position of its end effector to match the location of the user’s hand. To further enhance the naturalness of the robot’s motion, we employed inverse kinematics to calculate the joint transformations to make the robot move in synchrony with the user’s hand.
@@ 3.2. Path History
Trajectory visualization is a popular method to improve motion intent communication, as it allows users to preview the robot’s plan before execution. Previous research has shown that visualizing proposed robot motion using a mixed reality head-mounted display led to a 16% increase in accuracy and a 62% decrease in task completion time compared to a 2D display when labeling robot arm motions as either colliding or non-colliding with objects (Rosen et al., 2017). Inspired by these findings, we incorporated a virtual line trajectory using Unity’s line renderer component. The system displayed a straight green line between the robot’s end effector coordinate and the user’s hand coordinate to minimize the occlusion of the AR robot and objects. By utilizing the Jacobian matrix to relate joint velocities to end effector velocity, applying a damped least squares approach to compute joint angle updates, and iteratively updating the joint angles, our inverse kinematics implementation enables the end effector to move in approximately straight-line motions towards the desired goal pose. Thus, linear interpolation sufficiently approximates the end effector’s actual trajectory. To support real-time control and flexibility, we included a redo button that reverts the robot to the last saved waypoint and removes the corresponding trajectory. We opted for a redo functionality instead of projecting the robot’s goal pose before movement because our entire system is AR-based, ensuring safety and efficiency when moving the robot directly to the goal pose.

### v3

@@ Abstract.
The increasing affordability of robot hardware is accelerating the integration of robots into everyday activities. However, training a robot to automate a task requires expensive trajectory data where a trained human annotator moves a physical robot to train it. Consequently, only those with access to robots produce demonstrations to train robots. In this work, we remove this restriction with EVE, an iOS app that enables everyday users to train robots using intuitive augmented reality visualizations, without needing a physical robot. With EVE, users can collect demonstrations by specifying waypoints with their hands, visually inspecting the environment for obstacles, modifying existing waypoints, and verifying collected trajectories. In a user study ($N=14$, $D=30$) consisting of three common tabletop tasks, EVE outperformed three state-of-the-art interfaces in success rate and was comparable to kinesthetic teaching—physically moving a physical robot—in completion time, usability, motion intent communication, enjoyment, and preference ($mean_{p}=0.30$). EVE allows users to train robots for personalized tasks, such as sorting desk supplies, organizing ingredients, or setting up board games. We conclude by enumerating limitations and design considerations for future AR-based demonstration collection systems for robotics.
augmented reality, robotics, demonstration collection
Equal Contribution.
journalyear: 2024copyright: rightsretainedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676413isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Mixed / augmented realityccs: Computer systems organization Robotics
@@ 1. Introduction
Figure 1. EVE allows everyday users to collect data to train a real robot using intuitive augmented reality (AR) visualizations without needing a physical robot. Our application enables users to move the AR robot by setting waypoints with hand gestures, visually inspect the real-world environment to avoid obstacles, modify the collected trajectory, and verify the data collection by replaying the task with the AR robot. Videos of AR visualizations, features, and real-world deployment are available at https://junwang0510.github.io/EVE/.
[…]
An example collection with EVE. A person is holding an iPad with one hand and using the other to rearrange items on the table (left). Four examples of AR visualizations that assist the user in gathering data are shown (middle left). A neural network diagram represents using the collected data to train a real robot (middle right). Three images demonstrate the real-world deployment (right).
@@
The decreasing costs of robot components are accelerating the integration of robots into everyday life. With companies offering robotic arms at a few hundred dollars, portable, lightweight, and mobile robots are being trained to assist individuals with a wide range of everyday tasks (Elephant Robotics, 2023; Robot, 2023; Shaw et al., 2023a). However, training a robot to automate a task often requires expensive manually curated human trajectories. Annotators are hired and trained to kinesthetically move or teleoperate a physical robot. The robot’s joints and gripper are moved into a sequence of configurations to accomplish a task. For example, to train a robot to clean a table, annotators move the gripper towards a cloth, close the gripper to grasp the cloth, and then begin moving it around to wipe the table.
@@
There are no available systems for everyday users to train robots without first purchasing an expensive robot. Users who are considering purchasing a robot to automate a task in their homes need to first purchase the robot. Next, they need to learn how to move bulky machinery to produce useful trajectories. Consequently, only a limited number of roboticists and developers with access to robots will determine the range of consumer robotic applications.
@@
Recently, a data collection mechanism, called AR2-D2, demonstrated the possibility of training robots on simple tasks without needing a physical robot (Duan et al., 2023). AR2-D2 converts videos of people manipulating objects in their homes into robot trajectories. AR2-D2 uses computer vision models to automatically capture the 3D structure of the environment for training robot policies. AR2-D2 demonstrated that the quality of data gathered using videos is comparable with data collected using traditional kinesthetic or teleoperated approaches with real robots. However, AR2-D2 was only evaluated on simple, short-horizon tasks such as pressing a computer mouse or picking up a plastic bowl. Ecologically valid user tasks, such as cleaning a table, require multiple interactions between the user and their environment, resulting in longer videos. Since the tasks are longer, users are more likely to take wrong actions and need to backtrack and redo particular steps. Thus, while AR2-D2 shows that training a robot might not require a real robot, there is a need to explore the possible designs that make such a system usable and ecologically valid from a user-centric perspective.
We introduce EVE, an iOS app that allows everyday users to train robots using an augmented reality (AR) robot, without requiring a physical robot. To train a real robot, users use EVE to project an AR robot in their environment and guide its behavior. Users can specify waypoints that define the trajectory the AR robot should follow. During collection, users can visually inspect the movement of the AR robot from different angles to ensure that the AR robot avoids obstacles. Users can modify the collected trajectory by reverting the AR robot’s position to a previous waypoint. EVE also contains a number of visualization features that provide users with confidence that their trajectory could be used to train a real robot.
@@
To understand the opportunities and challenges of AR-based trajectory collection, we conduct a formative study with $10$ participants with varying experience in robotics, comparing AR2-D2 to the initial prototype of EVE with three different AR visualizations. From this study, we identify six usability challenges users faced and implemented seven additional features to improve EVE’s accuracy, flexibility, and verifiability.
@@
We design a user study ($N=14$) to quantitatively compare EVE against state-of-the-art trajectory collection interfaces on three tabletop tasks. EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in task completion time, system usability, motion intent communication, user enjoyment, and overall preference ($mean_{p}=0.30$). Participants customized their workflows with multiple visualizations and features to collect demonstrations for complex, long-horizon tasks. Additionally, we train machine learning policies with the collected trajectories and deploy the policies on a real robot for the toggle switch task. Specifically, we collect $6$ trajectories using both AR2-D2 and EVE and train policies for $30,000$ iterations. We find that on $30$ evaluations, policies trained with EVE resulted in $10\%$ higher accuracy.
EVE facilitates personalized robot training in everyday environments, allowing users to collect demonstrations tailored to their specific needs and adapt to various robot platforms and motion planners. There are still numerous unanswered questions: the current system lacks support for bimanual manipulation and mobile robot data collection, which are crucial for many real-world tasks like washing dishes and cleaning floors. Future research should investigate the use of AR head-mounted displays for bimanual control and the integration of enhanced depth cues for more precise manipulation. Moreover, incorporating Simultaneous Localization and Mapping (SLAM) techniques with EVE could enable multi-view data collection, which is essential for mobile robots. Enhancements in task variability are also necessary, such as through more efficient motion planners and hardware improvements, to ensure robust and accurate performance in dynamic and unstructured settings.
@@ 2. Related work
As we aim to enhance the accuracy, usability, and verifiability of an AR-based trajectory collection interface for robotics, our work builds upon prior research on human trajectory collection mechanisms and the use of AR to assist in motion intent communication.
2.1. Trajectory collection for robotics
A common approach for collecting demonstration data includes kinesthetic teaching, where an annotator physically guides the robot through a desired trajectory (Argall et al., 2009). Despite the direct haptic feedback, this mechanism can be slow and tedious (Osentoski et al., 2010) as users have to physically move a bulky, heavy machine by rotating its joints into desired configurations.
Another popular collection method is teleoperation, where the user controls the robot remotely using specialized controllers, such as 3D space mouse (Zhu et al., 2022), game or virtual reality (VR) controllers (Ebert et al., 2021; Brohan et al., 2022; Seo et al., 2023; Mahadevan et al., 2022), smartphones (Mandlekar et al., 2018), and haptic devices (Shaw et al., 2023b; Toedtheide et al., 2023). These specialized devices are costly and challenging to use due to high latency and unintuitive controls for modifying multiple robot joints (Akgün and Subramanian, 2011).
More recent interfaces capture human expert demonstrations with physical robots using specialized hardware like ALOHA (Zhao et al., 2023), GELLO (Wu et al., 2023), DexCap (Wang et al., 2024), and UMI (Chi et al., 2024). These hardware resemble gloves in the shape of the robot gripper, offering users greater flexibility in manipulating objects with the robot without needing to control the entire robot arm. However, these studies often prioritize data collection efficiency and robot performance, with minimal user studies examining user preference and usability (Kapadia et al., 2017; Szafir and Szafir, 2021).
@@ 2.1. Trajectory collection for robotics
Prior work compares kinesthetic teaching and teleoperation on simple tasks such as manipulating cups (Fischer et al., 2016; Akgün and Subramanian, 2011). The common finding is that kinesthetic teaching is preferred over teleoperation based on collection time, ease of use, and accuracy of demonstrations. By contrast, our paper proposes using an iOS app with AR to train robots. We focus on three realistic tabletop tasks with varying levels of difficulty, leading to a more comprehensive comparison between different collection interfaces.
Most closely related to our work is AR2-D2 (Duan et al., 2023), a data collection mechanism that does not require real robots or trained human annotators, and resituates trajectory collection outside a robotics lab with AR. AR2-D2 shows that users require no training in contrast to common demonstration collection interfaces. However, AR2-D2’s utility is limited; it focuses on simple, short-horizon tasks like clicking a button. They also do not explore the design space for developing a prototype for everyday users.
Our system, EVE, focuses on the human-centric aspect of the trajectory collection process with new visualizations and new features derived from a formative study. It allows users to select suitable collection methods based on different ecologically valid tasks. Furthermore, we examine the system’s accuracy, speed, usability, and verifiability through three common, long-horizon tabletop tasks.
@@ 2.2. Augmented reality and robotics
Traditional approaches to improving human-robot interaction, such as through robots’ movements (Dragan et al., 2013; Mainprice et al., 2010), gestural motion (Huang and Mutlu, 2013; Chidambaram et al., 2012), and gaze outputs (Mutlu et al., 2006; Kalegina et al., 2018), are limited due to the challenge of presenting expressive physical feedback that goes beyond internal capabilities. By contrast, AR provides visual feedback in one’s line of sight, opening up exciting new opportunities for human-robot interaction research. A recent taxonomy of AR and robotics has identified five purposes and benefits of visual augmentation (Suzuki et al., 2022). We focus on analyzing work related to motion intent communication since our goal is to improve system usability by allowing users to easily control and understand the goal of the AR robot.
@@
AR interfaces help communicate the robot’s intention to the user through spatial information. Prior work has experimented with different designs based on factors such as path, waypoints, velocity, arrival times, and shadows (Pascher et al., 2023; Walker et al., 2018; Zollmann et al., 2014; Chen et al., 2021). Most relevant to our work is Rosen et al. (Rosen et al., 2017), who visualized the robotic arm’s motion intent by overlaying planned robot poses with a mixed reality head-mounted display. Evaluating against 2D displays and no visualizations, they found that the AR visualization allows users to more quickly ($-62\%$ in time) and accurately ($+16\%$ in accuracy) identify where the robot is going to move. They also iterated on the system to allow the robot to replan an intended trajectory with MoveIt motion planner (mov, 2023) with the same start and end points, with the final trajectory chosen by the user (Rosen et al., 2019). This design has been shown to improve task completion time, precision, accuracy, usability, and task load compared to the baseline systems. Similarly, EVE uses an AR robot to communicate how a real robot might move with the collected waypoints. We also allow users to reselect the trajectory they desire by modifying waypoints. Unlike prior works, EVE is an iOS app, and the AR visualization is applied with a phone or a tablet. This form factor enables more people to collect demonstrations for training robots compared to wearable displays. Furthermore, instead of projecting the entire pose of the robot in AR, we move the AR robot directly to the goal pose due to the lack of safety concerns caused by physical collisions.
3. EVE initial prototype
@@
Figure 2. The initial prototype of EVE includes three new AR visualizations. AR kinesthetic teaching allows the robot to track the user’s hand movements in real-time. Path history displays the trajectory along collected waypoints, enabling users to revert to previous waypoints. Invisible robot allows users toggle off the robot’s body, showing a cylinder that represents the robot’s end effector position and gripper state.
[…]
Three new AR visualizations in the formative study: AR Kinesthetic Teaching (top row), Path History (middle row), and Invisible Robot (bottom row). Each row consists of four images illustrating using the respective visualization to push the cube.
@@
While prior work has compared common demonstration collection interfaces and AR visualizations, there is a lack of research investigating user preferences for AR-based demonstration collection systems. Informed by the usability issues of AR2-D2, we developed the initial EVE prototype with three new AR visualizations: AR kinesthetic teaching, path history, and invisible robot (Figure 2). Before we describe these visualizations, we first describe the user workflow for collecting trajectories.
3.1. Trajectory collection in AR
Our system enables users to collect AR trajectories for training real robots. After selecting a task for data collection, the user projects an AR robot into the real-world environment using the iOS app and guides the robot by specifying trajectories through hand movements. The application captures the trajectory of the AR robot by saving RGB-D data and waypoint coordinates. For instance, to train the robot to push a drawer, the user opens the app and directs the camera to the target location for the AR robot. By tapping the screen, they designate where the AR robot will appear. The user can then move the gripper by defining waypoints with their hands. The system captures and records RGB-D data along with waypoint coordinates during this interaction, saving the data to the device’s album for subsequent use in training the real robot policy.
While AR2-D2 demonstrated the possibility of using a similar workflow to collect robot trajectories, we identified several usability issues. Firstly, users must specify waypoints before pressing a button to move the AR robot. This can be unintuitive as users may struggle to visualize the robot’s configuration at the waypoint. Besides, users cannot modify the collected waypoints, making it difficult to gather data for complex, long-horizon tasks without restarting the application. Lastly, when the AR robot is placed near the target object, it can occupy a significant portion of the device’s screen, further complicating the data collection process. These usability challenges remained invisible in AR2-D2 because of their focus on simple, short-horizon tasks, such as picking up a plastic bowl and pressing a computer mouse.
@@ 3.2. AR kinesthetic teaching (AR-KT)
AR-KT enhances system usability by synchronizing the robot’s movements with the user’s hand. To ensure natural robot motion, the system tracks the user’s hand position in real-time and leverages inverse kinematics to calculate the joint transformations necessary for robot movement. Prior work highlights that kinesthetic teaching generally surpasses teleoperation in system usability, collection time, and task success rate (Akgün and Subramanian, 2011), primarily due to the intuitiveness and immediate feedback of manually guiding the robot. Moreover, recent research (Maric et al., 2024) has shown that human demonstration using a virtual marker excels in quality, time, and workload compared to kinesthetic teaching when drawing patterns. AR-KT is similar to this approach as it allows users to move their hands freely without the need to manually adjust robot joint configurations to match desired positions.
@@ 3.3. Path history
Path history leverages Unity’s Line Renderer component to display virtual trajectory lines between collected waypoint coordinates. Our inverse kinematics implementation uses the Jacobian matrix to relate joint velocities to end effector velocity and applies a damped least squares approach to compute joint angle updates. This iterative adjustment allows the end effector to move in near-straight-line paths toward the target pose (Figure 2). Trajectory visualization is a well-established technique for enhancing motion intent communication, allowing users to preview a robot’s planned path before execution. Prior research indicates that AR-based trajectory displays improve intent clarity and decrease task completion times (Walker et al., 2018; Rosen et al., 2017, 2019).
To facilitate real-time adjustments and flexibility, we implement a redo button that reverts the robot to its last saved waypoint and clears the associated trajectory. We choose this functionality over projecting the robot’s goal pose before movement because our system is entirely AR-based, ensuring both safety and efficiency when moving the robot directly to the goal pose.